How many grams of sodium are needed to produce 3.90 moles of sodium oxide
____Na + ____O2 —-> ____Na2O

(both 2 are supposed to be smaller)
1
Expert's answer
2019-05-22T07:03:35-0400

4Na + O2 = 2Na2O

n(Na) = 3.9*4 / 2

n(Na) = 7.8 mole

m(Na) = 7.8*23 = 179.4 g